CoreLinux++ FAQ

The Corelinux Consortium

Revision: 1.2 Created on May 8, 2000

Revised on July 28, 2000

Abstract

This document is the Corelinux++ Frequently Asked Question guide. It should answer any of the questions you asked about Corelinux++ and its development process.

Contents

1 CoreLinux++ and Open Source 2 1.1 Are you really Open Source? . . 2 1.2 What are you offering? . . . . . . 2 1.3 Will the class libraries and frame-

works be portable? . . . . . . . . 2 1.4 What about the C++ Standards? 2 1.5 How are you funded? . . . . . . . 2

2 Contacting the CoreLinux++ Consortium 2 2.1 Web Resources . . . . . . . . . . 2

2.1.1 Does CoreLinux++ have a

web page? . . . . . . . . . 2 2.1.2 What about a project page? 2

2.2 Internet Resources . . . . . . . . 3

2.2.1 Is there a mailing list? . . 3 2.3 Defect Reporting or Suggestions 3

2.3.1 How do I report a defect or make a suggestion? . . 3

3 General Information 3

4 File and source code access 3

4.1 Is there an ftp site to get the cur-

rent tarball? . . . . . . . . . . . . 3 4.2 Is the source tree available? … 3

4.3 Is there another way? . . . . . . 4

5 Documentation 4 5.1 C++ Coding Style . . . . . . . . 4 5.1.1 Is there a document de-

scribing this? . . . . . . . 4

5.2 Object Oriented Analysis and Design . . . . . . . . . . . . . . . . 4 5.2.1 Is there a document de-

scribing the process? … 4

5.3 Developers . . . . . . . . . . . . . 4 5.3.1 What do I need to know

to build the CoreLinux++

class libraries? . . . . . . 4 5.3.2 I want to be involved, how

do I do this? . . . . . . . 4 5.3.3 I want to contribute but I

don’t want to join. Can I

do this? . . . . . . . . . . 4 5.3.4 What do I need to know to

use the CoreLinux++ bi-

naries? . . . . . . . . . . . 4

5.4 Class Library Documentation . . 4 5.4.1 Is there documentation of

the class libraries? …. 4 5.4.2 How is the documentation

created? . . . . . . . . . . 4 2 CONTACTING THE CORELINUX++ CONSORTIUM

Copyright notice

CoreLinux++ Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999, 2000 CoreLinux Consortium
Revision:  1.2 Last Modified:  July 28, 2000

This material may be distributed only subject to the terms and conditions set forth in the Open Publication License.

1 CoreLinux++ and Open Source

1.1 Are you really Open Source?

Yes, all of the documentation and graphics are licensed under the Open Publication License, and all of the source code under the Free Software Foundations Lesser General Public License (LGPL).

1.2 What are you offering?

At a minimum, a coherent set of standards for Object Oriented Analysis and Design with C++ implentations in Linux. Ultimately we will be providing a set of class libraries that provide other developers with common classes, patterns, framework, and service abstractions.

1.3 Will the class libraries and frameworks be portable?

Our only goal is to be portable across Linux platforms (distributions and hardware). Because of this we can achieve native performance in the class implementations and avoid the weight gain when trying to be all things to all platforms.

1.4 What about the C++ Standards?

We encourage and exploit the standards as set in the ISO 14882 (1998E) C++ Standard.

1.5 How are you funded?

Free time, the belief in Open Source, Linux, other organizations’ belief in the same, and sweat.

2 Contacting the CoreLinux++ Consortium

2.1 Web Resources

2.1.1 Does CoreLinux++ have a web page?

The CoreLinux++ web site.

2.1.2 What about a project page?

The CoreLinux++ Project Page has links to file downloads, message forums, mailing lists, defect reporting and general information.

4 FILE AND SOURCE CODE ACCESS  2.2 Internet Resources

2.2 Internet Resources

2.2.1 Is there a mailing list?

Go to the Subscription page and follow the directions.

2.3 Defect Reporting or Suggestions

2.3.1 How do I report a defect or make a suggestion?

Sign on to the Project Page and select the Bug Report link.

2.3.2: What do the defect groups mean?

We have identified a number of groups for reporting defects, or making suggestions:

Features Something you would like to see from CoreLinux++ that isn’t in plan or already a feature request

Severity 1 Defect core dump, corrupt data, etc. The REALLY BAD STUFF

Severity 2 Defect incorrect IO, breaks existing applications, etc

Severity 3 Defect Class interfaces or implementation not robust enough, compiler errors, etc.

Defects General catch all for other defects not characterized by above.

3 General Information

4 File and source code access

4.1 Is there an ftp site to get the current tarball?

Anonymous logon to

ftp://corelinux.sourceforge.net

follow to

/pub/corelinux/

4.2 Is the source tree available?

The project’s CVS tree can be checked out through anonymous (pserver) CVS with the following instruction set. The module you wish to check out must be specified as the modulename. When prompted for a password for anonymous, simply press the Enter key.

   cvs -d:pserver:anonymous@cvs.corelinux.sourceforge.net:/cvsroot/corelinux login
cvs -d:pserver:anonymous@cvs.corelinux.sourceforge.net:/cvsroot/corelinux co modulename

Updates from within the module’s directory do not need the -d parameter.

5 DOCUMENTATION  4.3 Is there another way?

4.3 Is there another way?

As an alternative, the CoreLinux++ CVS tree is archived daily

5 Documentation

5.1 C++ Coding Style

5.1.1 Is there a document describing this? Refer to the Coding Standards page, see [Consortium, 2000a].

5.2 Object Oriented Analysis and Design

5.2.1 Is there a document describing the process? Refer to the Analysis and Design page, see [Consortium, 2000b]..

5.3 Developers

5.3.1 What do I need to know to build the CoreLinux++ class libraries? Read the Developers Guide or just run make from the downloaded release /corelinux directory.

5.3.2 I want to be involved, how do I do this? Join the mailing list, read the CoreLinux++ process documentation, and if you are still interested send e-mail to me

5.3.3 I want to contribute but I don’t want to join. Can I do this? Yes. Use our projects Patch manager facility. We can’t promise that it will be incorporated, but it will be reviewed and you will be contacted.

5.3.4 What do I need to know to use the CoreLinux++ binaries? Examine the examples in /corelinux/src/testdrivers for a start, see following section

5.4 Class Library Documentation

5.4.1 Is there documentation of the class libraries? Under construction by CoreLinux++ Class Reference.

5.4.2 How is the documentation created? We are juggling between two systems that generate HTML from comments in the source headers. Luckily they both work off the same style of commenting.

REFERENCES  REFERENCES

References

The Corelinux Consortium. The Corelinux C++ Coding Standards. The Corelinux Consortium, 1.3 edition, May 2000a. http://corelinux.sourceforge.net/cppstnd.php.

The Corelinux Consortium. The Corelinux Object Oriented Design Standards. The Corelinux Consortium, 1.3 edition, May 2000b.